OpenStudy (anonymous):

find the slope and y-intercept of the equation of the line. -3y +6x=15

OpenStudy (anonymous):

You need to solve for y first... what is your first step to solve for y?

OpenStudy (anonymous):

-3y +6x=15 y=(-6x+15)/(-3) y=2x-5 find y intercept plug o for x y=2(0)-5 y=-5

OpenStudy (anonymous):

Once you have it in y = form the number in front of the x is the slope (2) the number added/subtracted to/from the x is the y-intercept (-5)

OpenStudy (anonymous):

y=mx+b is the equation of a line so once you solve the equation for y then just look at the coefficient of x (which is m for the slope) and if + a number comes next then the y-intercept is positive, if - a number, y-int is negative. m=slope, b=y-int